A valószínűségszámítás gyakorlaton való részvétel

dátum:  IX. 11. IX. 18. IX. 25. X. 2. X. 9. X.16. X.30

diákok:

Békési Zita  +  +  +  −  +  +
Bóka Zsolt  +  +  −  +  +  +
Éliás Anikó  +  −  +  −  −  −
Fábián Andrea  +  +  −  −  −  +
Fink Evelin  +  −  −  +  +  +
Gyöngy Imre  +  −  −  −  −  −
Garai Krisztina  +  +  +  +  +  −
Hadházi István  −  −  −  +  −  −
Herka Erzsébet  +  +  −  −  −  +
Homoki István  +  +  +  −  −  +
Juhász Anikó  −  +  +  +  −  +
Kocsis Zsuzsanna  +  +  +  +  +  +
László Anikó  −  +  −  +  −  −
Mezei Tamás  +  −  −  +  −  −
Miló Viktor  +  −  −  −  +  −
Magyar Tünde  +  +  +  +  +  +
Molnár Renáta  +  −  +  +  +  +
Nagy Anikó  +  +  +  +  +  +
Nagy Gábor  +  −  +  +  +  −
Nagy Mónika  +  −  +  +  +  +
Nagyházi Melinda  −  +  +  −  +  +
Oláh Erika  −  −  +  +  −  +
Pelei Csaba  −  +  +  −  +  −
Ragány Szilvia  +  +  +  +  +  +
Sebestyén Hajnalka  +  +  +  +  −  +
Szilvási Krisztina  +  +  −  +  +  +
Sztankó Patricia  +  +  +  +  +  +
Szuetta Judit  −  −  −  +  −  −
Tóth Viktória  −  +  −  −  +  +
Wittner Ferenc  +  −  −  −  −  −
Zsitnyár Judit  +  +  +  +  +  +

dátum:  XI 6. XI. 13. XI. 27. XII. 4. XII. 11.

diákok:

Békési Zita  +  −  −
Bóka Zsolt  −  +  −
Éliás Anikó  −  −  −
Fábián Andrea  −  −  −
Fink Evelin  −  −  −
Gyöngy Imre  −  −  −
Garai Krisztina  −  +  +
Hadházi István  −  −  −
Herka Erzsébet  −  +  −
Homoki István  +  −  +
Juhász Anikó  −  +  +
Kocsis Zsuzsanna  +  +  +
László Anikó  +  −  −
Mezei Tamás  −  −  −
Miló Viktor  −  −  −
Magyar Tünde  +  +  +
Molnár Renáta  +  +  +
Nagy Anikó  +  +  +
Nagy Gábor  +  +  +
Nagy Mónika  +  −  −
Nagyházi Melinda  −  −  +
Oláh Erika  +  +  −
Pelei Csaba  −  −  −
Ragány Szilvia  +  −  +
Sebestyén Hajnalka  +  +  +
Szilvási Krisztina  +  +  +
Sztankó Patricia  +  +  +
Szuetta Judit  −  −  −
Tóth Viktória  +  +  −
Wittner Ferenc  −  −  +
Zsitnyár Judit  +  +  +